OpenStudy (anonymous):

can somebody check my answer? Write an equation of a line passing through the points (-3, 6) and (-5, 3). Write the final answer in slope-intercept form?

OpenStudy (anonymous):

m=y2 -y1 x2 -x1 3-(6) = -3 -5-(-3) -2 y- y1=m(x-x1) y+6=3/2(x-3) y+6=3/2x-9/2 y+6-6=3/2x-9/2-6 y=3/2x-21/2

OpenStudy (anonymous):

m=y2 -y1 x2 -x1 3-(6) = -3 -5-(-3) -2 y- y1=m(x-x1) y+6=-3/2(x-3) y+6=-3/2x-9/2 y+6-6=-3/2x+9/2-6 y=-3/2x-3/2

pooja195 (pooja195):

\(\huge\color{pink}{WELCOME~TO~OPENSTUDY!!!}\) Your answers look correct well done :D

OpenStudy (anonymous):

thank you very much!

pooja195 (pooja195):

haha your welcome :)